# MoPD: Mixture-of-Prompts Distillation for Vision-Language Models

### v1

@@ IV-B Results on Few-shot Classification
In the few-shot classification setting, each method is trained on all the classes of a dataset with few-shot labeled samples and evaluated on the rest data of this dataset. Following [11], we perform few-shot classification with different numbers of shots, specifically 1, 2, 4, 8, and 16 shots. The performance of all the methods on the 11 datasets is shown in Fig. 3. According to the results, MoPD achieves the best average performance among all baseline methods. Those results further demonstrate the good generalization ability of the proposed MoPD method when the samples of downstream tasks are limited.
@@ IV-C Results on Domain Generalization
According to results shown in Table II, we can observe that MoPD outperforms all baseline methods except ProGrad in the source domain, and MoPD performs better than all baseline methods on ImageNet-Sketch and ImageNet-A. In terms of average performance on all the four target domains, MoPD performs the best among all baseline methods, which confirms that the soft prompts learned in MoPD are more domain-generalizable.
TABLE II: Performance of various methods under the domain generalization setting. All methods use 16-shot samples from each of the 1000 classes on ImageNet except CLIP. The best and second-best results are marked in bold and underlined, respectively.
@@
IV-D Ablation Study
We conduct the ablation study to demonstrate the effect of mixture-of-prompts distillation loss $\zeta_{\mathrm{MPD}}$, mixture-of-prompts selection loss $\zeta_{\mathrm{MPS}}$, and the selection of gating network under the base-to-new generalization setting.
The results are listed in Table III, where the baseline is CoOp, SiPD built on CoOp is formulated in Section III-B, MoPD-R denotes a variant of MoPD by randomly selecting teacher prompts instead of using the gating network, and MoPD (w/o $\zeta_{\mathrm{MPS}}$) denotes MoPD without minimizing $\zeta_{\mathrm{MPS}}$.
@@ IV-D Ablation Study
TABLE III: Ablation study for the mixture-of-prompts distillation loss $\zeta_{\mathrm{MPD}}$, mixture-of-prompts selection loss $\zeta_{\mathrm{MPS}}$, and gating network.
@@ IV-E Parameter Analysis
IV-F Further Analysis
To investigate the robustness of MoPD to noisy prompts, we add noisy prompts into the hard prompt pool and compare MoPD with MoPD-R under the base-to-new generalization setting on the UCF101 and DTD datasets, where noisy prompts are randomly generated based on grammatically plausible sentence templates with incoherent words generated by ChatGPT [42, 43], such as “In the whispering forest, a [CLASS] dreams of stars” (more examples can be found in Supplementary Material).
According to the results shown in Table IV, when the hard prompt pool is full of noisy prompts (i.e., 24N and 12N), the performance of both MoPD and MoPD-R is clearly worse than MoPD and MoPD-R with 12T. Nevertheless, even in situations where the hard prompt pool is exclusively composed of noisy prompts, MoPD is capable of discerning relatively superior teacher prompts, resulting in higher harmonic mean accuracy compared to MoPD-R. When the hard prompt pool is a mixture of noisy prompts and task-related prompts, the higher the proportion of noisy prompts in the prompt pool, the worse the performance (in terms of base, new, and harmonic mean accuracy) of MoPD-R. However, the performance of MoPD with 12T shows almost no difference when compared to MoPD with 12T+12N and MoPD with 12T+24N. This is because the gating network can help select suitable teacher prompts. Those experimental results demonstrate the robustness of MoPD against noisy prompts.
@@ IV-F Further Analysis
TABLE IV: Robustness of MoPD to noisy prompts under the base-to-new generalization setting on the UCF101 and DTD datasets. ‘N’ denotes noisy prompts, ‘T’ denotes task-related prompts, and ‘+’ denotes the mixture of the two types of prompts.
@@ Acknowledgments
This work is supported by National Key R&D Program of China (No. 2022ZD0160300) and NSFC key grant 62136005.
Appendix A Implementation Details
A-A Hard Prompt Pool
The hard prompt pool consisting of task-related hard prompts is simply designed through synonym substitution and sentence transformation on prompt templates of CLIP [1] and CoOp [11]. Unless otherwise specified, both MoPD and MoPD-R methods utilize the following hard prompt pool in the experiments. SiPD performs prompt distillation by employing the first prompt from the hard prompt pool.
For ImageNet, Caltech101, StanfordCars, and SUN397, the hard prompt pool is
‘‘a photo of a [CLASS].’’
‘‘a picture of a [CLASS].’’
‘‘an image of a [CLASS].’’
‘‘a photo, classified as [CLASS].’’
‘‘a photo, categorized as [CLASS].’’
‘‘a photo, labeled as [CLASS].’’
‘‘a picture, classified as [CLASS].’’
‘‘a picture, categorized as [CLASS].’’
‘‘a picture, labeled as [CLASS].’’
‘‘an image, classified as [CLASS].’’
‘‘an image, categorized as [CLASS].’’
‘‘an image, labeled as [CLASS].’’
For Food101, the hard prompt pool is
‘‘a photo of [CLASS], a type of food.’’
‘‘a picture of [CLASS], a type of food.’’
‘‘an image of [CLASS], a type of food.’’
‘‘a photo of food, classified as [CLASS].’’
‘‘a photo of food, categorized as
  [CLASS].’’
‘‘a photo of food, labeled as [CLASS].’’
‘‘a picture of food, classified as
  [CLASS].’’
‘‘a picture of food, categorized as
  [CLASS].’’
‘‘a picture of food, labeled as [CLASS].’’
‘‘an image of food, classified as
  [CLASS].’’
‘‘an image of food, categorized as
  [CLASS].’’
‘‘an image of food, labeled as [CLASS].’’
For OxfordPets, Flowers102, and FGVCAircraft, the hard prompt pool is
‘‘a photo of a [CLASS], a type of {}.’’
‘‘a picture of a [CLASS], a type of {}.’’
‘‘an image of a [CLASS], a type of {}.’’
‘‘a photo of a {}, classified as [CLASS].’’
‘‘a photo of a {}, categorized as
  [CLASS].’’
‘‘a photo of a {}, labeled as [CLASS].’’
‘‘a picture of a {}, classified as
  [CLASS].’’
‘‘a picture of a {}, categorized as
  [CLASS].’’
‘‘a picture of a {}, labeled as [CLASS].’’
‘‘an image of a {}, classified as
  [CLASS].’’
‘‘an image of a {}, categorized as
  [CLASS].’’
‘‘an image of a {}, labeled as [CLASS].’’
For DTD, the hard prompt pool is
‘‘a photo of [CLASS], a type of texture.’’
‘‘a picture of [CLASS], a type of
  texture.’’
‘‘an image of [CLASS], a type of texture.’’
‘‘a photo of a texture, classified
  as [CLASS].’’
‘‘a photo of a texture, categorized
  as [CLASS].’’
‘‘a photo of a texture, labeled
  as [CLASS].’’
‘‘a picture of a texture, classified
  as [CLASS].’’
‘‘a picture of a texture, categorized
  as [CLASS].’’
‘‘a picture of a texture, labeled
  as [CLASS].’’
‘‘an image of a texture, classified
  as [CLASS].’’
‘‘an image of a texture, categorized
  as [CLASS].’’
‘‘an image of a texture, labeled
  as [CLASS].’’
For EuroSAT, the hard prompt pool is
‘‘a centered satellite photo of
  [CLASS].’’
‘‘a centered satellite picture of
  [CLASS].’’
‘‘a centered satellite image of
  [CLASS].’’
‘‘a centered satellite photo,
  classified as [CLASS].’’
‘‘a centered satellite photo,
  categorized as [CLASS].’’
‘‘a centered satellite photo,
  labeled as [CLASS].’’
‘‘a centered satellite picture,
  classified as [CLASS].’’
‘‘a centered satellite picture,
  categorized as [CLASS].’’
‘‘a centered satellite picture,
  labeled as [CLASS].’’
‘‘a centered satellite image,
  classified as [CLASS].’’
‘‘a centered satellite image,
  categorized as [CLASS].’’
‘‘a centered satellite image,
  labeled as [CLASS].’’
For UCF101, the hard prompt pool is
‘‘a photo of a person doing [CLASS].’’
‘‘a picture of a person doing [CLASS].’’
‘‘an image of a person doing [CLASS].’’
‘‘a photo of a person doing, classified
  as [CLASS].’’
‘‘a photo of a person doing, categorized
  as [CLASS].’’
‘‘a photo of a person doing, labeled
  as [CLASS].’’
‘‘a picture of a person doing, classified
  as [CLASS].’’
‘‘a picture of a person doing, categorized
  as [CLASS].’’
‘‘a picture of a person doing, labeled
  as [CLASS].’’
‘‘an image of a person doing, classified
  as [CLASS].’’
‘‘an image of a person doing, categorized
  as [CLASS].’’
‘‘an image of a person doing, labeled
  as [CLASS].’’
[…]
A-B Noisy Hard Prompt Pool
The term ”noisy prompts” refers to prompts that are unrelated to the downstream task. Taking inspiration from the generation of random prompts by [44], we employ a similar approach to generate noisy prompts. These noisy prompts are random, grammatically plausible sentence templates with incoherent words generated by ChatGPT [42, 43]. We prompt the language model with the input: “Please generate 12 random prompt templates by sampling grammatically plausible random sentences that contain incoherent words, with length between 5 and 20 words, such as “a photo of a [CLASS]”, where [CLASS] is the class name.” The noisy prompt pool of size 24 is shown as follow.
‘‘A collection of [CLASS] floating in the
  sky.’’
‘‘[CLASS] wrapped in a mysterious bubble.’’
‘‘An unexpected encounter with a [CLASS].’’
‘‘A [CLASS] hidden among the stars.’’
‘‘A vibrant dance of the [CLASS].’’
‘‘The laughter of a [CLASS] echoing through
  the forest.’’
‘‘A [CLASS] emerging from the depths of the
  ocean.’’
‘‘A journey through the pages of a
  [CLASS].’’
‘‘The whispers of a [CLASS] in the wind.’’
‘‘A [CLASS] perched on the edge of a
  cliff.’’
‘‘A [CLASS] illuminated by the moonlight.’’
‘‘The song of a [CLASS] resonating in the
  mountains.’’
‘‘Under the luminous sky, a [CLASS] dances
  with shadows.’’
‘‘In the whispering forest, a [CLASS]
  dreams of stars.’’
‘‘A [CLASS] floats in a teacup filled
  with stormy clouds.’’
‘‘With a melody of forgotten songs, a
  [CLASS] paints the sunrise.’’
‘‘A [CLASS] juggles moonbeams at the edge
  of twilight.’’
‘‘Inside a clockwork, a [CLASS] unravels
  the fabric of time.’’
‘‘Beneath the rainbow waterfall, a [CLASS]
  plays a silent tune.’’
‘‘A [CLASS] weaves dreams into the fabric
  of the cosmos.’’
‘‘In the heart of an ancient book, a
  [CLASS] discovers hidden worlds.’’
‘‘A [CLASS] dances on the petals of an
  upside-down flower.’’
‘‘Amongst the whispers of the wind, a
  [CLASS] listens to the earth’s
  heartbeat.’’
‘‘A [CLASS] rides the waves of a sunbeam
  across the ocean of eternity.’’
Note that in Table IV, the notation “12N” denotes that MoPD and MoPD-R are trained utilizing the first 12 noisy prompts listed in the table. While the notation ”24N” denotes training with all 24 noisy prompts in the table.
Appendix B Experimental Details
B-A Details for Datasets
The detailed statistics of the 11 datasets for the base-to-new generalization setting are shown in Table A5, where the soft prompts are learned with 16 instances per base class.
TABLE A5: The detailed statistics of datasets under the base-to-new generalization setting.
Datasets Number of classes Number of samples
Base New Train Test-Base Test-New
ImageNet 500 500 8000 16402 16300
Caltech101 50 50 800 1549 916
OxfordPets 19 18 304 1881 1788
StanfordCars 98 98 1568 4002 4039
Flowers102 51 51 816 1053 1410
Food101 51 50 816 15300 15000
FGVCAircraft 50 50 800 1666 1667
SUN397 199 198 3184 9950 9900
DTD 24 23 384 864 828
EuroSAT 5 5 80 4200 3900
UCF101 51 50 816 1934 1849
Appendix C Additional Results on Load Balance
Inspired by SMoP [45], we adopt load balance for exploring diverse teacher prompts by adding a scaled Gaussian noise $\delta\sim\mathcal{N}\left(0,1\right)$, which decreases within the training process, onto $(\boldsymbol{f}\cdot\boldsymbol{W}_{g})$ in Eq. (5) in Sec. III-B. Consequently, Eq. (5) is modified as
$G(\boldsymbol{f})=\mathrm{Softmax}(\mathrm{KeepTop}((\boldsymbol{f}\cdot\boldsymbol{W}_{g}\cdot(\mathbf{I}+\delta\cdot p))),T))$ (10)
where $\mathbf{I}$ is the identity matrix, and $p\in\left[0,1\right]$ is the ratio of current to maximum iterations.
The results for the base-to-new generalization setting is shown in Table A6, where MoPD-L represents a variant of MoPD that uses load balance, the task-related hard prompt pool means that its hard prompts are simply designed by synonym substitution and sentence transformation on CLIP’s prompt templates, and the mixed hard prompt pool means the mixture of task-related and task-unrelated hard prompt pool.
TABLE A6: The effect of load balance under the base-to-new generalization setting on all datasets. The numbers in red and blue denote the best performance from different prompt pools, respectively.
Dataset Set Task-related hard prompt pool Mixed hard prompt pool
MoPD MoPD MoPD MoPD MoPD MoPD
-R -L -R -L
Average Base 81.39 81.28 81.40 80.74 81.09 81.04
New 73.52 74.64 74.69 73.92 74.76 74.52
H 77.26 77.82 77.90 77.18 77.80 77.64
ImageNet Base 76.87 76.83 76.87 76.97 77.00 76.93
New 70.10 70.10 70.43 70.17 69.87 70.30
H 73.33 73.31 73.51 73.41 73.26 73.47
Caltech 101 Base 98.07 97.90 98.07 97.90 97.83 97.97
New 94.50 95.50 94.90 95.27 94.73 94.57
H 96.25 96.69 96.46 96.57 96.26 96.24
Oxford Pets Base 95.13 94.80 95.17 95.13 95.30 95.33
New 97.40 97.17 97.47 97.27 96.83 97.57
H 96.25 95.97 96.31 96.19 96.06 96.44
Stanford Cars Base 75.70 75.27 75.63 75.33 75.50 75.83
New 72.70 73.57 73.10 70.33 72.53 72.40
H 74.17 74.41 74.34 72.74 73.99 74.08
Flowers 102 Base 96.30 96.13 95.80 95.63 96.00 95.43
New 72.90 73.27 74.43 72.03 73.53 74.23
H 82.98 83.16 83.77 82.17 83.28 83.51
Food101 Base 90.60 90.80 90.77 90.57 90.73 90.70
New 91.60 91.67 91.67 91.73 91.97 91.97
H 91.10 91.23 91.22 91.15 91.35 91.33
FGVC Aircraft Base 36.50 36.47 37.43 36.93 35.03 37.90
New 34.70 36.50 35.13 34.73 35.83 35.07
H 35.58 36.48 36.24 35.80 35.43 36.43
SUN397 Base 80.87 81.27 81.13 80.87 80.77 80.90
New 76.17 76.57 76.47 76.00 76.83 76.43
H 78.45 78.85 78.73 78.36 78.75 78.60
DTD Base 78.27 78.43 77.27 77.57 79.00 75.13
New 55.57 56.60 57.47 57.33 57.80 59.33
H 64.99 65.75 65.92 65.93 66.76 66.30
EuroSAT Base 85.40 84.83 85.77 80.07 83.80 83.77
New 65.37 71.73 72.50 71.00 74.80 69.33
H 74.05 77.73 78.58 75.26 79.04 75.87
UCF101 Base 81.60 81.37 81.53 81.13 81.00 81.53
New 77.70 78.37 78.07 77.27 77.67 78.50
H 79.60 79.84 79.76 79.15 79.30 79.99
With the task-related hard prompt pool, compared to MoPD-R and MoPD-L, MoPD achieves the best average performance (in terms of base, new, and harmonic mean accuracy). MoPD-L is slightly inferior to MoPD but also achieves good performance. With the mixed hard prompt pool, MoPD-L achieves slightly better performance than MoPD. Due to the negligible difference in average performance between MoPD and MoPD-L, we use the MoPD method in all the experiments.

### v2

@@ IV-B Results on Few-shot Classification
In the few-shot classification setting, each method is trained on all the classes of a dataset with few-shot labeled samples and evaluated on the rest data of this dataset. Following [11], we perform few-shot classification with different numbers of shots, specifically 1, 2, 4, 8, and 16 shots. The performance of all the methods on the 11 datasets is shown in Fig. 3. According to the results, MoPD achieves the best average performance among all TP and HP baseline methods. While MaPLe and MetaPrompt outperform the TP methods in terms of the average performance under 8 and 16 shots, their effectiveness diminishes under 1, 2, and 4 shots, achieving only comparable performance to TP methods. This limitation likely arises from insufficient parameter optimization under extreme data scarcity. In contrast, MoPD-MMP exhibits better average performance than all baseline methods under all shots. Those results further demonstrate the good generalization ability of the proposed MoPD and MoPD-MMP methods even when the samples of downstream tasks are limited.
@@ IV-C Results on Domain Generalization
According to results shown in Table II, we can observe that MoPD and MoPD-MMP outperform all baseline methods except ProGrad and MetaPrompt in the source domain. In terms of the average performance on all four target domains, MoPD and MoPD-MMP outperform all baseline methods and especially MoPD-MMP possesses the highest overall average accuracy of 60.63%, which confirms that the soft prompts learned in MoPD are more domain-generalizable.
TABLE II: Performance of various methods under the domain generalization setting. All methods use 16-shot samples from each of the 1000 classes on ImageNet except CLIP. The best and second-best results are marked in bold and underlined.
@@
MaPLe 70.72 64.07 49.15 50.90 76.98 60.28
MetaPrompt 72.60 65.50 49.30 49.03 77.20 60.28
MoPD-MMP 71.93 65.40 49.30 50.70 77.10 60.63
IV-D Results on Cross-Dataset Evaluation
In the cross-dataset evaluation setting that evaluates whether prompts learned from a source domain can generalize to unseen target domains, prompts are trained using 16-shot samples from each of the 1000 classes on ImageNet and then evaluated on the other 10 datasets. According to results shown in Table III, the proposed MoPD and MoPD-MMP methods perform comparably to the SOTA method MaPLe while using significantly fewer tunable parameters (8K vs. 52K vs. 3.55M), and demonstrate competitive performance compared to other baseline methods.
TABLE III: Performance of various methods under the cross-dataset evaluation setting. All methods use 16-shot samples from each of the 1000 classes on ImageNet except CLIP. The best results are marked in bold. The numbers in the parentheses following each method name indicate the number of tunable parameters.
Source Target
ImageNet Caltech101 OxfordPets StanfordCars Flowers102 Food101 FGVCAircraft SUN397 DTD EuroSAT UCF101 Avg.
CoOp (2K) 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88
CoCoOp (35K) 71.02 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21 65.74
ProGrad (2K) 72.24 91.52 89.64 62.39 67.87 85.40 20.61 62.47 39.42 43.46 64.29 62.71
KgCoOp (2K) 70.66 93.92 89.83 65.41 70.01 86.36 22.51 66.16 46.35 46.04 68.50 65.51
MoPD (8K) 71.77 94.20 90.50 66.30 71.60 86.80 23.90 67.00 47.10 43.80 69.60 66.08
MaPLe (3.55M) 70.72 93.53 90.49 65.57 72.23 86.20 24.74 67.01 46.49 48.06 68.69 66.30
MetaPrompt (61K) 72.60 93.83 89.87 65.13 70.83 85.63 24.50 67.10 46.73 47.17 66.47 65.73
MoPD-MMP (52K) 71.93 94.30 90.70 65.70 71.10 86.30 23.40 66.90 46.40 48.30 68.90 66.20
IV-E Ablation Study
We conduct the ablation study to demonstrate the effect of mixture-of-prompts distillation loss $\zeta_{\mathrm{MPD}}$, mixture-of-prompts selection loss $\zeta_{\mathrm{MPS}}$, and the selection of gating network under the base-to-new generalization setting. The results are listed in Table IV, where the baseline is CoOp, SiPD built on CoOp is formulated in Section III-B, MoPD-R denotes a variant of MoPD by randomly selecting teacher prompts instead of using the gating network, and MoPD (w/o $\zeta_{\mathrm{MPS}}$) denotes MoPD without minimizing $\zeta_{\mathrm{MPS}}$.
@@ IV-E Ablation Study
TABLE IV: Ablation study for the mixture-of-prompts distillation loss $\zeta_{\mathrm{MPD}}$, mixture-of-prompts selection loss $\zeta_{\mathrm{MPS}}$, and gating network.
@@
Effect of knowledge transfer methods. We conduct ablation studies on knowledge transfer methods. We evaluated four distinct methods: 1) KL divergence-based knowledge distillation used in MoPD, which minimizes the KL divergence between prediction probability distributions of hard and soft prompts; 2) MMD-based knowledge distillation, which minimizes the Maximum Mean Discrepancy [51] between prediction probability distributions of hard and soft prompts; 3) Cosine similarity-based knowledge transfer [14], which maximizes cosine similarity between textual embeddings of hard and soft prompts; and 4) $\ell_{1}$ loss-based knowledge transfer [28], which minimizes the $\ell_{1}$ loss between textual embeddings of hard and soft prompts. According to results shown in Table V, we can see that the proposed MoPD method that is with KL divergence-based knowledge distillation outperforms alternative knowledge transfer methods in terms of average harmonic mean accuracy, demonstrating that the proposed knowledge distillation method in MoPD is effective.
TABLE V: Effect of different knowledge transfer methods on the average performance of all 11 datasets under the base-to-new generalization setting.
Method Base New H
MoPD-$\ell_{1}$ 81.74 73.16 77.21
MoPD-cos 80.24 74.46 77.24
MoPD-MMD 80.63 73.18 76.72
MoPD (ours) 81.40 74.69 77.90
Computational Cost. The computational cost of MoPD is shown in Table VI. We can observe that MoPD introduces negligibly computational overhead compared to SiPD, while MoPD achieves performance gains over SiPD. Those results demonstrate that the proposed MoPD is time-efficient.
TABLE VI: Computational cost of MoPD, where the model is trained on the base classes of ImageNet. $H$ denotes the pool’s total number of hard prompts, ‘ms’ denotes the millisecond per image, and ‘H (Avg.)’ denotes the average harmonic mean accuracy of 11 datasets.
Method Learnable para. Train.time Infer.time Memory H (Avg.)
SiPD 4K 31min36s 2.33ms 9044MB 77.04
MoPD (4+$H$)K +15s +0ms +5.8MB 77.90
@@ IV-F Parameter Analysis
IV-G Further Analysis
Robustness of MoPD. To investigate the robustness of MoPD to noisy prompts, we add noisy prompts into the hard prompt pool and compare MoPD with MoPD-R under the base-to-new generalization setting on the UCF101 and DTD datasets, where noisy prompts are randomly generated based on grammatically plausible sentence templates with incoherent words generated by ChatGPT [52, 53], such as “In the whispering forest, a [CLASS] dreams of stars” (more examples can be found in Supplementary Material).
According to the results shown in Table VII, when the hard prompt pool is full of noisy prompts (i.e., 24N and 12N), the performance of both MoPD and MoPD-R is clearly worse than MoPD and MoPD-R with 12T. Nevertheless, even in situations where the hard prompt pool is exclusively composed of noisy prompts, MoPD is capable of discerning relatively superior teacher prompts, resulting in higher harmonic mean accuracy compared to MoPD-R. When the hard prompt pool is a mixture of noisy prompts and task-related prompts, the higher the proportion of noisy prompts in the prompt pool, the worse the performance (in terms of base, new, and harmonic mean accuracy) of MoPD-R. However, the performance of MoPD with 12T shows almost no difference when compared to MoPD with 12T+12N and MoPD with 12T+24N. This is because the gating network can help select suitable teacher prompts. Those experimental results demonstrate the robustness of MoPD against noisy prompts.
@@ IV-G Further Analysis
Visualization. We use t-SNE [54] to analyze prediction distributions on new classes of the EuroSAT dataset under the 16-shot base-class training setting. As shown in Fig. 5, we compare CLIP, CoOp, CoCoOp, ProGrad, KgCoOp, and MoPD, where the numbers 0 to 4 represent the categories of Pasture Land, Permanent Crop Land, Residential Buildings, River, and Sea or Lake, respectively. The visualization shows that KgCoOp’s prediction manifold closely aligns with CLIP’s zero-shot distribution. CoOp, CoCoOp, and ProGrad exhibit larger intra-class variance and ambiguous inter-class boundaries. Notably, MoPD achieves better inter-class separation, particularly for similar category pairs like Pasture Land (dark blue) and Permanent Crop Land (green), River (grey) and Sea or Lake (light blue). Those visualization results indicate the enhanced discriminative capability of the proposed MoPD method. In summary, those visualizations corroborate the good generalization of MoPD on new classes.
(a) CLIP
(b) CoOp
(c) CoCoOp
(d) ProGrad
(e) KgCoOp
(f) MoPD
Figure 5: Visualization of different prediction distributions via t-SNE.
Analysis on Failure Case. We observe that MoPD underperforms CoCoOp and ProGrad on the OxfordPets dataset under the base-to-new generalization setting. This can be attributed to small category shift between the downstream task and CLIP’s pretraining data, as noted in [55]. Consequently, MoPD extracts limited generalization benefits from hard prompts, leading to inferior results for new classes and harmonic mean accuracy in this specific dataset.
TABLE VII: Robustness of MoPD to noisy prompts under the base-to-new generalization setting on the UCF101 and DTD datasets. ‘N’ denotes noisy prompts, ‘T’ denotes task-related prompts, and ‘+’ denotes the mixture of the two types of prompts.
@@ Acknowledgments
This work is supported by National Key R&D Program of China (No. 2022ZD0160300) and NSFC grant (No. 62136005).